OpenStudy (anonymous):

A rock is dropped off the edge of a cliff and its distance (in feet) from the top of the cliff after t seconds is s(t)=16t^2. Assume the distance from the top of the cliff to the water below is 1024ft. When will the rock strike the water?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you would solve 16t^2 = 1024 for t

jimthompson5910 (jim_thompson5910):

what do you get

OpenStudy (anonymous):

so... 16(1024)^2...

jimthompson5910 (jim_thompson5910):

no you're solving 16t^2 = 1024 for t not plugging in t = 1024

OpenStudy (anonymous):

O...8

jimthompson5910 (jim_thompson5910):

yep, it will take 8 seconds

OpenStudy (anonymous):

So how do i find the average velocity when the time interval is [7,8]?

jimthompson5910 (jim_thompson5910):

use the following AROC = (f(x2) - f(x1))/(x2 - x1)

jimthompson5910 (jim_thompson5910):

x1 = 7 x2 = 8

OpenStudy (anonymous):

can u do a walkthrough?

jimthompson5910 (jim_thompson5910):

s(7) = ???

jimthompson5910 (jim_thompson5910):

s(t)=16t^2 s(7)=16(7)^2 s(7) = ???

OpenStudy (anonymous):

49

jimthompson5910 (jim_thompson5910):

16(7)^2 = 16(49) = 784

jimthompson5910 (jim_thompson5910):

s(7) = 784

jimthompson5910 (jim_thompson5910):

what is s(8)

OpenStudy (anonymous):

1024

OpenStudy (anonymous):

av=240?

jimthompson5910 (jim_thompson5910):

AROC = (s(x2) - s(x1))/(x2 - x1) AROC = (s(8) - s(7))/(8 - 7) AROC = (1024 - 784)/(8 - 7) AROC = (240)/(1) AROC = 240 so you got it

OpenStudy (anonymous):

Thanks so much for your help tonight, hope your not tired of me :)

jimthompson5910 (jim_thompson5910):

you're welcome and no I like helping, so you're fine